Shift Summary for day shift, 3 January 2000


Subject: Shift Summary for day shift, 3 January 2000
From: Vivian O'Dell (odell@fnal.gov)
Date: Wed Jan 12 2000 - 16:29:23 CST


              Shift Summary for day shift, 3 January 2000

On Shift: Julie, Patty

Hours of beam delivered at all intensities: 1.0
Hours of beam delivered at nominal intensity: 0.0
Hours of E799 data (special runs not included): 0.0
Hours of special runs (laser scans, muon runs...): 0.5

Accounting of lost time: Beam startup, tape drive allocation problems,
                           fpedsyn problems

========================================================================
Runs ended this shift:

Run 15366 Pedestal Run
Run 15369 BEAM_MU_4E12

========================================================================
Accesses during shift:

Ed and RickK for BA, mask, regenerator
Jason, Leo, Sydney to fix ET power strip
Julie and Jason H to swap divider for channel 983
BobT and RickK for BA timing study

========================================================================
Overview of shift:

   Setup for KAMI test, remade hv1440.exe and fpedsyn

   RickC and Herman make an access into NM2 to plug the WEST beam hole.
   East beam has 18" of Be (shadow absorber). Common absorber can be
   controlled by epicure.

   RUN PLAN details in LOGBOOK XXIII p. 169.

========================================================================
Description of Shift:

run spill time comment
- - ------------------------------------------------------------------------
              11:00 Regenerator pushed into east beam; Mask swung into
                        place.

15366 13:40 Pedestal run; cspd and adgn constants updated

              14:00 Rick Coleman calls to say that access to NM2 is
                        complete

              14:30 Beam arrives

15369 14:33 Muon Run
        
              14:35 DaveN fills the alcohol bubbler and increased flow
                        to DCs from 0.5 to 0.7 cfh

              14:50 Install DC alignment constants DCBM 15369

              15:00 Set common absorber to 20".

              15:30 Call MCR to ask them to remove beam stop.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

check to see if fpedsyn runs properly



This archive was generated by hypermail 2b27 : Wed Jan 12 2000 - 16:29:31 CST